INAF LBC+MODS queue observing, 2012 March 26-27

Observer: K. Boutsia, R. Speziali, M. Faccini
Telescope Operator: S. Allanson
Support Astronomer: O. Kuhn
Instrument: MODS + LBC RED

Summary

LBC is out of service due to a fault in the B science pc. Almost clear sky with some high cirrus

Data Summary

Project Objects FilterSorted descending ExpTime Notes
GILLI SDSSJ1411+1217 r 1.2hr seeing <1"
FIORE COSMOS3   4hr seeing <1"
DONNARUMMA Abell1703   1.2hr seeing <1"

Details

02:00 Collimating (seeing around 1")

02:15 starting the acquisition of Feige34
  • acqMODS Feige34.acq mods1r.20120327.0003-0005

02:23 Observation of Feige34
  • execMODS Feige34.obs mods1r.0006-0007/mods1b.0003

02:37 Acquisition of FIORE Cosmos3
  • acqMODS mods.55.FIORE3.acq mods1r.0008-0010
  • modsAlign mods.55.FIORE3.mms mods1r.20120327.0008.fits mods1r.20120327.0009.fits
   dX:   0.099 arcsec
   dY:  -1.597 arcsec
  dPA: -0.1914 degrees

02:55 Sci field FIORE Cosmos3
  • execMODS mods.55.FIORE3.obs mods1r.0011/mods1b.0004
  • 400 sec exposure and then abort due to misaligment of the targets in the slits

02:58 offsetting by hand by DX=0.369"

03:09 Through slit 60 sec image mods1r.0012

03:12 execMODS mods.55.FIORE3.obs mods1r.0013-0020/mods1b.0005-0012

07:30 Acquisition of DONNARUMMA A1703
  • acqMODS mods.201150.newmi2e.acq mods1r.0021-0023
  • modsAlign mods.201150.newmi2e.mms mods1r.20120327.0021.fits mods1r.20120327.0022.fits
   dX:   0.137 arcsec
   dY:  -0.740 arcsec
  dPA: -0.1130 degrees

07:53 Sci field DONNARUMMA A1703
  • execMODS mods.201150.newmi2e.obs mods1r.0024-0025

09:14 script abort after the second exposition

09:25 starting the acquisition of Feige66
  • acqMODS Feige66.acq mods1r.0026-0028
   dX:  -0.123 arcsec
   dY:  -0.246 arcsec

09:32 starting the observation of Feige66
  • execMODS Feige66.obs mods1r.0029-0030 / mods1b.0013
  • manual exp of 80 sec with red grating since 120 sec it's nearly saturated. mods1r.0031

09:49 SWITCH TO LBC

10:06 Gilli focusOB4

10:11 dofpia, /FIRST, /x2, /redonly

10:21 ri4-1 - fac

11:00 dofpia, /x2, /redonly

11:15 12 iterations to converge

11:16 ri4-1 - fac

11:50 dofpia, /x2, /redonly

12:08 13 iterations to converge

12:10 ri4-1 - fac

12:20 aborted cause high sky background

12:35 r sloan flat

12:50 telescope parked

12:54 calibration script for MODS:
  • execMODS -u grflat_8.cal mods1r 0032-0046 / mods1b 0014-0023 This set was requested to us by OSURC
  • execMODS -u grmosflats.55.FIORE3.cal mods1r 0047-0056 / mods1b 0024 - 0033
  • execMODS -u grmosflats.55.FIORE3new.cal mods1r 0057-0061 / mods1b 0034-0038
  • execMODS -u grmosflatsDonnaRm2.cal mods1r 0062-0071
  • execMODS -u SlitFlatLBT5.0.cal mods1r 0072-0076 / mods1b 0039-0043
  • execMODS -u grpixflats.cal mods1r 0077-0081 / mods1b 0044-0053
  • execMODS -u allbias.cal mods1r 0082-0091 / mods1b 0054-0063

03:08 End of night
  • execMODS ~MODSeng/modsScripts/Support/modsSleep.pro

  • Average seeing during MODS observations:
    20120327.jpg

-- RobertoSpeziali - 27 Mar 2012
Topic revision: r1 - 31 Jul 2024, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback